# HybridVLA: Collaborative Diffusion and Autoregression in a Unified Vision-Language-Action Model

### v2

@@ 3.2 Collaborative Training Recipe
Directly combining diffusion and autoregressive pose prediction within a single LLM presents challenges such as instability and inconsistencies in next-token prediction. Therefore, we propose a collaborative training recipe that encompasses token sequence formulation design, hybrid objectives, and structured training stages.
Token sequence formulation design. As shown in the upper part of Figure 2, the input token sequence during training comprises not only visual and language tokens but also robot state, diffusion noise, and autoregressive action tokens. For the robot state, we integrate it into the LLM to enhance temporal consistency in action generation. Instead of discretizing the robot state and merging it with the language query [56] (Type 3 of Table 1), we employ a learnable MLP to map the robot state directly into the word embedding space, $f_{r}\in\mathbb{R}^{B\times 1\times 4096}$. The motivation is that diffusion action tokens are generated during the subsequent next-token prediction process, using all preceding tokens as conditions. Introducing discrete robot states could negatively impact the diffusion prediction of continuous actions. For diffusion actions, we predict them through a diffusion denoising process, conditioned on the previous tokens. The denoising step $i$ and noisy actions $a^{i}_{t}$ are projected into the LLM’s word embeddings through an MLP, represented as continuous vectors. To seamlessly connect previous multi-modal tokens, diffusion tokens, and subsequent discrete tokens within a single sequence, we introduce special beginning-of-diffusion ($<$BOD$>$) and end-of-diffusion ($<$EOD$>$) tokens to encapsulate the diffusion tokens. This design not only clarifies the boundaries between the diffusion and autoregressive generation but also prevents confusion in the next-token prediction process, such as diffusion tokens directly predicting masked discrete tokens (Type 2 of Table 1). For autoregressive actions, we quantize the end-effector pose into discrete bins and replace part of the vocabulary in the LLM [47], which is then tokenized into a sequence of discrete tokens. Owing to the inherent nature of next-token prediction, both the question and the answer (discrete action GT) are fed into the LLM during training, while only the question is provided during inference. If the autoregressive prediction is placed before the diffusion tokens, it would lead to action GT leakage, serving as a condition for diffusion training (Type 4 of Table 1). Consequently, we place the diffusion token at the front explicitly providing continuous knowledge for the subsequent autoregressive generation.
Type1(Ours) Type2 Type3 Type4
0.66 0.56 0.61 0.57
@@
Table 1: Exploring and validating our proposed token sequence formulation. The model is trained using both diffusion and autoregressive generation but tested only on diffusion-based actions (HybridVLA-dif) across 10 simulation tasks.
Hybrid objectives. To simultaneously train diffusion and autoregressive action generation, we require two distinct loss functions. For diffusion part, following previous diffusion policies [14], we minimize the mean squared error between the predicted noise ($\epsilon_{\pi}$) from the VLA model and the GT noise ($\epsilon$). The loss function is defined as follows:
$L_{dif}=E_{a,i,c}||\epsilon-\epsilon_{\pi}(a^{i}_{t},i,c)||^{2}$
Where $\epsilon\sim N(0,1)$ and $c$ denotes the condition. Additionally, classifier-free guidance [30] is not used in order to ensure stable robot arm behavior [61]. For the autoregressive part, the cross-entropy loss ($L_{ce}$) is adopted to supervise the discrete output. With our designed token sequence formulation, the two losses can be seamlessly combined for collaborative penalization, defined as:
$L_{hybrid}=L_{dif}+L_{ce}$
@@
Structured training stage. After loading the pretrained VLM parameters, HybridVLA undergoes two training stages with hybrid objectives: large-scale pretraining on open-source robotic data and fine-tuning on self-collected simulation and real-world data. During pretraining, we train HybridVLA for 5 epochs on 35 datasets from Open X-Embodiment [69], DROID [46], ROBOMIND [101], and others. The pretrain datasets contain 760k robot trajectories, comprising 33m frames. Due to dataset differences, pretraining relies solely on single 2D observations, whereas fine-tuning relies on either single or multi-view observations, depending on the downstream task. The details of the pretraining dataset are provided in Appendix A.1.
@@ 3.3 Collaborative Action Ensemble
During inference, given visual, language, and robot state inputs, HybridVLA concurrently generates actions using both diffusion and autoregressive methods, and then ensembles the actions to produce more stable execution.
@@
Autoregressive actions. As shown in Figure 2, the autoregressive generation begins after the special token $<$EOD$>$. Similar to [47, 56], the generation of 7-DoF or 14-DoF actions closely mirrors the text generation process in an LLM. Unlike previous autoregressive VLA methods, HybridVLA’s autoregressive generation additionally conditions on the inherently continuous nature of actions from diffusion tokens, outperforming independent autoregressive approaches, as demonstrated in the ablation study.
@@
Diffusion actions. When generating diffusion actions, we append the special token $<$BOD$>$ after the previous condition tokens to indicate that the model should perform the denoising process. We employ DDIM [89] sampling with $n$ sampling steps. In HybridVLA, we find that integrating diffusion into the next-token prediction process not only improves action precision by fully leveraging the reasoning abilities and pretrained knowledge of VLMs but also mitigates performance degradation when reducing the number of inference denoising steps (e.g., $n=4$), as demonstrated in the ablation study. To accelerate the sampling process, we introduce the KV cache before the diffusion tokens, forwarding conditional information, the denoising timestep, and pure noise only during the initial sampling step. In subsequent steps, the cached keys and values from the first pass are reused, while only the timestep and noise are iteratively forwarded. This strategy eliminates redundant computations and improves inference speed.
Models box laptop lid seat down to dustpan fridge on base out off hanger rack plants S.R.$\uparrow$ speed
ManipLLM (7B) [56] 0.50 0.80 0.40 0.20 0.80 0.35 0.10 0.25 0.15 0.20 0.38 2.2 Hz
OpenVLA (7B) [47] 0.65 0.40 0.75 0.60 0.80 0.20 0.35 0.15 0.10 0.10 0.41 6.3 Hz
$\pi_{0}$ (2.6B) [8] 0.90 0.60 1.00 0.30 0.90 0.25 0.35 0.75 0.05 0.45 0.55 13.8 Hz
CogACT (7B) [52] 0.80 0.85 0.90 0.65 0.90 0.50 0.60 0.35 0.25 0.25 0.60 9.8 Hz
HybridVLA-dif (7B) 0.85 0.75 1.00 0.80 0.95 0.50 0.50 0.30 0.70 0.25 0.66 9.4 Hz
HybridVLA (7B) 0.85 0.95 1.00 0.90 1.00 0.50 0.50 0.70 0.50 0.50 0.74 6.1 Hz
HybridVLA (2.7B) 1.00 0.80 0.90 0.80 0.90 0.25 0.20 0.45 0.25 0.25 0.58 12.3 Hz
Table 2: Comparison of our proposed method and baselines on RLBench. We train all methods in the Multi-task setting [87] and report the success rates (S.R.) for each task. The success condition follows the definition in RLBench. “HybridVLA-dif” refers to action prediction relying solely on the diffusion process. (7B), (2.7B), and (2.6B) refer to the sizes of the LLM used in the VLA model.
@@
Ensembled actions. After obtaining the two types of actions under our collaborative training recipe, we empirically observe two phenomena. 1) Different action types demonstrate varying performance across tasks. Diffusion-based predictions excel in precise manipulation tasks, such as Phone on base and Close laptop lid, while autoregressive predictions perform better in tasks requiring scene semantic reasoning, such as Water plants and Frame off hanger. 2) The confidence of autoregressive action tokens serves as a reliable indicator of action quality. In over 80% of successfully completed test samples, the average confidence of autoregressive action tokens exceeds 0.96. Quantitative evaluations are provided in Appendix B.1 and B.2. Therefore, as shown at the end of Figure 2, we use the mean confidence of autoregressive tokens ($c^{ar}_{t+1}$) to guide the action ensemble. If the confidence exceeds $\theta$ ($\theta=0.96$), we consider the autoregressive action ($a^{ar}_{t+1}$) sufficiently accurate and perform an average operation with the diffusion action ($a^{d}_{t+1}$). Otherwise, we rely solely on the diffusion action to control the robot. Additionally, to accelerate inference, HybridVLA-dif relies solely on diffusion-based action generation while still collaboratively learning both action generation types during training to enhance mutual reinforcement.
@@ 4.1 Simulation Experiment
Simulation benchmark. To systematically evaluate, we select the RLBench [36] benchmark in the CoppeliaSim simulator, which contains 10 different tabletop tasks. These tasks, performed using a Franka Panda robot and a front-view camera, include Close box, Close Laptop, Toilet seat down, Sweep to dustpan, Close fridge, Phone on base, Take umbrella out, Frame off hanger, Wine at rack, and Water plants. The data are collected using pre-defined waypoints and the Open Motion Planning Library [90]. Following the frame-sampling method used in previous works [87, 24, 40], we construct the training dataset, with each task consisting of 100 trajectories.
@@
Training and Evaluation Details. We compare our method against four previous SOTA VLA models, including autoregressive-based approaches such as ManipLLM [56] and OpenVLA [47], as well as diffusion-based methods like $\pi_{0}$[8] and CogAct[52] with a DiT-base action head. Meanwhile, we categorize our method into three modes: HybridVLA (7B), HybridVLA (2.7B), and HybridVLA-dif (7B). All modes are jointly trained using our proposed collaborative training recipe; however, HybridVLA-dif relies solely on diffusion-based action generation during inference. To ensure a fair comparison, we load the official pretrained parameters provided by each method, adhering to their respective training settings. For HybridVLA, the single-view RGB input is resized to $224\times 224$, and the robot state is consistent with predicted actions (7-DOF end-effector poses). During training, we use the AdamW optimizer with a constant learning rate of 2e-5. As shown in Figure 2, we only update the LLM and injected MLP parameters. Our models are trained for 300 epochs on 8 NVIDIA A800 GPUs with mixed-precision training. For evaluation, following [47, 52], all methods are tested with 20 rollouts from the latest epoch.
@@
Quantitative Results. As shown in Table 2, HybridVLA(7B) achieves an average success rate of 74% across 10 distinct tasks, outperforming the previous SOTA autoregressive-based VLA (OpenVLA) and diffusion-based VLA (CogACT) by 33% and 14%, respectively. These results demonstrate that our method seamlessly introduces diffusion modeling within the LLM and effectively combines the advantages of two types of pose prediction. Remarkably, compared to CogACT and $\pi_{0}$, HybridVLA-dif also achieves performance improvements of 6% and 11%, respectively. These results highlight that, unlike previous approaches that attach the diffusion head after the VLM, our method effectively leverages the inherent capabilities of the VLM to unlock the full potential of diffusion-based action generation within the VLA model. Finally, HybridVLA(2.7B) delivers satisfactory results, confirming our method’s effectiveness in enhancing VLM manipulation capabilities across different model sizes.
Inference Speed. By harnessing VLM’s reasoning capabilities and conducting large-scale robotic pretraining, our method reduces denoising steps to 4 without any performance degradation, as shown in Figure 3. In Table 2, when tested on an NVIDIA 4090D GPU, HybridVLA-dif (7B) and HybridVLA (2.7B) achieve satisfactory frequencies comparable to CogACT (7B) and $\pi_{0}$ (2.6B), thanks to the reduced denoising steps and efficient KV cache of our method. All models are loaded with bfloat16 precision for inference. Notably, while KV cache has been used in previous autoregressive VLA methods [47, 56], we first integrate it into the LLM’s diffusion-based action prediction.

### v3

@@ 3.2 Collaborative Training Recipe
Combining continuous diffusion and discrete autoregressive action generation within a single LLM presents challenges such as instability and inconsistency in the next-token prediction process. To address this, we propose a collaborative training recipe that includes a token sequence formulation, hybrid objectives, and structured training stages.
Token sequence formulation design. As shown in Figure 2, this design aims to organize multimodal tokens, such as robot state, diffusion noise, and autoregressive tokens, within the LLM’s embedding space into a unified and ordered token sequence, enabling coordination between the two generation paradigms during the next-token prediction process. For the robot state, we integrate it into the LLM to enhance temporal consistency in action generation. Instead of discretizing the robot state and merging it with the language query li2024manipllm (Type 3 of Table 3.2), we employ a learnable MLP to map the robot state directly into the embedding space, $f_{r}\in\mathbb{R}^{B\times 1\times 4096}$.
@@
Table 1: Token sequence formulations. All models are trained with both generation methods. Dif and AR denote evaluations using only diffusion-generated or autoregressive-generated actions, respectively, across 10 RLBench tasks.
Paradigm Type1(Ours) Type2 Type3 Type4
Dif 0.66 0.56 0.61 0.57
AR 0.62 0.54 0.59 0.60
The motivation is that diffusion action tokens are generated using all preceding tokens as conditions. Introducing discrete robot states could negatively impact the diffusion prediction of continuous actions. For diffusion-based actions, we predict them through a diffusion denoising process. During training, the denoising step $i$ and noisy actions $a^{i}_{t}$ are projected into the LLM’s word embeddings through an MLP, represented as continuous vectors. To seamlessly connect previous multimodal tokens, diffusion tokens, and subsequent discrete tokens within a sequence, we introduce special beginning-of-diffusion ($<$BOD$>$) and end-of-diffusion ($<$EOD$>$) tokens to encapsulate the diffusion tokens. This design not only clarifies the boundaries between diffusion and autoregressive generation but also prevents confusion in the next-token prediction process, such as avoiding diffusion tokens directly predicting masked discrete tokens (Type 2 of Table 3.2). For autoregressive actions, we quantize the end-effector pose into discrete bins and replace part of the vocabulary in the LLM kim2024openvla , which is then tokenized into a sequence of discrete tokens. Due to the autoregressive nature of LLMs touvron2023llama2 , both the question and the answer, including the discrete action ground truth (GT), are provided during training, whereas only the question is available during inference. Therefore, placing autoregression before the diffusion tokens may cause action GT leakage (Type 4 in Table 3.2), as all preceding tokens (which contain GT during training) serve as conditions in diffusion modeling. To avoid this, we position diffusion tokens before autoregression to explicitly provide continuous latent conditions for subsequent token prediction. Moreover, since diffusion operates on noise, it naturally circumvents the risk of information leakage.
Hybrid objectives. To simultaneously train diffusion and autoregressive action generation, we require two distinct loss functions. For the diffusion part, following previous diffusion policies chi2023diffusion , we minimize the mean squared error between the predicted noise ($\epsilon_{\pi}$) from the VLA model and the GT noise ($\epsilon$). The loss function is defined as follows: $L_{dif}=E_{a,i,c}||\epsilon-\epsilon_{\pi}(a^{i}_{t},i,c)||^{2}$, where $\epsilon\sim N(0,1)$ and $c$ denote the condition. Additionally, classifier-free guidance ho2022classifier is not used in order to ensure stable robot arm behavior liu2024rdt . For the autoregressive part, the cross-entropy loss ($L_{ce}$) is adopted to supervise the discrete output. With our designed token sequence formulation, the two losses can be seamlessly combined for collaborative penalization, defined as: $L_{hybrid}=L_{dif}+L_{ce}$. Since $L_{dif}$ and $L_{ce}$ penalize a shared LLM backbone, their gradients are jointly backpropagated, allowing the model to effectively absorb both the continuous characteristics of diffusion-based actions and the semantic reasoning representations derived from autoregressive generation, thereby enabling mutual reinforcement between the two paradigms.
@@
Structured training stage. After loading the pretrained VLM parameters, HybridVLA undergoes two training stages with hybrid objectives: large-scale pretraining on open-source robotic data and fine-tuning on self-collected data. During pretraining, we train HybridVLA for 5 epochs on 35 datasets o2023open ; khazatsky2024droid ; khazatsky2024droid . The pretrain datasets contain 760k robot trajectories, comprising 33m frames. Due to dataset differences, pretraining relies solely on single 2D observations, whereas fine-tuning relies on either single or multi-view observations, depending on the downstream task. The details of the pretraining dataset are provided in Appendix A.1.
@@ 3.3 Collaborative Action Ensemble
During inference, HybridVLA takes visual, language, and robot state inputs to concurrently generate actions via both diffusion and autoregressive methods, and ensembles them for execution.
@@
Autoregressive actions. As shown in Figure 2, the autoregressive generation begins after the special token $<$EOD$>$. Unlike previous autoregressive VLA methods kim2024openvla ; li2024manipllm , HybridVLA’s autoregressive generation additionally conditions on continuous action representations derived from the latent features of diffusion tokens. This results in superior manipulation performance compared to independent autoregressive discrete generation paradigms that lack explicit continuous latent conditioning, as demonstrated in the ablation study.
@@
Diffusion actions. When generating diffusion actions, we append the special token $<$BOD$>$ after the previous condition tokens to indicate that the model should perform the denoising process. We employ DDIM song2020denoising with $n$ sampling steps. In HybridVLA, we observe that the number of inference denoising steps can be reduced to 4 without causing any performance degradation. As illustrated in the denoising process of Figure 2, we repeat the process for 4 DDIM steps by feeding the noisy sample from the previous step into the LLM to predict the noise token for the current step, thereby fully leveraging the LLM’s contextual reasoning capabilities. In this way, we effectively inherit the LLM’s pretrained knowledge and seamlessly integrate diffusion generation into the next-token prediction process. Moreover, since we deliberately place the diffusion action tokens before the autoregressive tokens, the autoregressive predictions cannot be directly used as diffusion conditions. However, as discussed in the previous section, both generation methods share the same LLM backbone, which is jointly trained with hybrid objectives. As a result, the LLM is able to absorb the unique knowledge from each generation paradigm, thereby enhancing its overall representation. To accelerate the sampling process, we introduce the KV cache before the diffusion tokens, forwarding conditional information, the denoising timestep, and pure noise only during the initial sampling step. In subsequent steps, the cached keys and values from the first pass are reused, while only the timestep and noise are iteratively forwarded. This strategy eliminates redundant computations and improves inference speed.
@@
Ensembled actions. After obtaining the two types of actions under our collaborative training recipe, we empirically observe two phenomena. 1) Different action types demonstrate varying performance across tasks. Diffusion-based predictions excel in precise manipulation tasks, such as Phone on base and Close laptop lid, while autoregressive predictions perform better in tasks requiring scene semantic reasoning, such as Water plants and Frame off hanger. 2) The confidence of autoregressive tokens serves as a reliable indicator of action quality. In over 80% of successfully completed test samples, the average confidence of autoregressive action tokens exceeds 0.96. Quantitative evaluations are provided in Appendix  B.1 and  B.2. Therefore, as shown in Figure 2, we use the mean confidence of autoregressive tokens ($c^{ar}_{t+1}$) to guide the action ensemble. If the confidence exceeds $\theta$ ($\theta=0.96$), we consider the autoregressive action ($a^{ar}_{t+1}$) sufficiently accurate and perform an average operation with the diffusion action ($a^{d}_{t+1}$). Otherwise, we rely solely on the diffusion action to control the robot.
@@ 4.1 Simulation Experiment
Table 2: Comparison of HybridVLA and baselines on RLBench. We train all methods in the Multi-task setting shridhar2022peract and report the success rates (S.R.). The success condition follows the definition in RLBench. (7B), (2.7B), and (2.6B) refer to the sizes of the LLM used in the VLA model.
Models box laptop lid seat down to dustpan fridge on base out off hanger rack plants S.R. & Var speed
ManipLLM (7B) li2024manipllm 0.50 0.80 0.40 0.20 0.80 0.35 0.10 0.25 0.15 0.20 0.38 ±$\pm 0.042$ 2.2 Hz
OpenVLA (7B) kim2024openvla 0.65 0.40 0.75 0.60 0.80 0.20 0.35 0.15 0.10 0.10 0.41 ±$\pm 0.038$ 6.3 Hz
$\pi_{0}$ (2.6B) black2024pi_0 0.90 0.60 1.00 0.30 0.90 0.25 0.35 0.75 0.05 0.45 0.55 ±$\pm 0.035$ 13.8 Hz
CogACT (7B) li2024cogact 0.80 0.85 0.90 0.65 0.90 0.50 0.60 0.35 0.25 0.25 0.60 ±$\pm 0.041$ 9.8 Hz
HybridVLA-dif (7B) 0.85 0.75 1.00 0.80 0.95 0.50 0.50 0.30 0.70 0.25 0.66 ±$\pm 0.040$ 9.4 Hz
HybridVLA (2.7B) 1.00 0.80 0.90 0.80 0.90 0.25 0.20 0.45 0.25 0.25 0.58 ±$\pm 0.031$ 12.3 Hz
HybridVLA (7B) 0.85 0.95 1.00 0.90 1.00 0.50 0.50 0.70 0.50 0.50 0.74 ±$\pm 0.037$ 6.1 Hz
@@
Simulation benchmark. To systematically evaluate, we select the RLBench james2020rlbench benchmark in the CoppeliaSim simulator, which contains 10 different tabletop tasks. These tasks, performed using a Franka Panda robot and a front-view camera, include Close box, Close Laptop, Toilet seat down, Sweep to dustpan, Close fridge, Phone on base, Take umbrella out, Frame off hanger, Wine at rack, and Water plants. The data are collected using pre-defined waypoints and the Open Motion Planning Library sucan2012open . Following the frame-sampling method used in previous works shridhar2022peract ; goyal2023rvt ; jia2024lift3d , we construct the training dataset, with each task consisting of 100 trajectories.
@@
Training and Evaluation Details. We compare our method with four previous SOTA VLA models, including autoregressive-based approaches such as ManipLLM li2024manipllm and OpenVLA kim2024openvla , as well as diffusion-based methods like $\pi_{0}$ black2024pi_0 and CogAct li2024cogact with a DiT-base action head. Meanwhile, we categorize our method into three modes: HybridVLA (7B), HybridVLA (2.7B), and HybridVLA-dif (7B). All modes are jointly trained using our proposed collaborative training recipe; however, HybridVLA-dif relies solely on diffusion-based action generation during inference. To ensure a fair comparison, we load the official pretrained parameters provided by each method, adhering to their respective training settings. For HybridVLA, the single-view RGB input is resized to $224\times 224$, and the robot state is consistent with predicted actions (7-DOF end-effector poses). During training, we use the AdamW optimizer with a fixed learning rate of 2e-5 to update both the LLM and the injected MLP parameters. Our models are trained for 300 epochs on 8 NVIDIA A800 GPUs with mixed-precision training. For evaluation, we follow kim2024openvla ; li2024cogact and test all methods using 20 rollouts from the latest epoch checkpoint. Since RLBench employs a sampling-based motion planner goyal2024rvt , we evaluate each model three times per task and report the mean success rate along with its variance.
@@
Quantitative Results. As shown in Table 2, HybridVLA(7B) achieves an average success rate of 74% across 10 distinct tasks, outperforming the previous SOTA autoregressive-based VLA (OpenVLA) and diffusion-based VLA (CogACT) by 33% and 14%, respectively. These results demonstrate that our method effectively combines the two generation approaches within a shared LLM backbone, simultaneously capturing the continuous characteristics of diffusion-based actions and the pretrained semantic reasoning capabilities learned through autoregression. Remarkably, compared to CogACT and $\pi_{0}$, HybridVLA-dif also achieves performance improvements of 6% and 11%, respectively. These results highlight that, unlike previous approaches which attach the diffusion head after the VLM, our method more effectively leverages the VLM’s pretrained knowledge to fully unlock the potential of diffusion prediction. Finally, HybridVLA(2.7B) delivers satisfactory results, confirming our method’s effectiveness in enhancing VLM manipulation capabilities across different model sizes. Inference Speed. In Table 2, when tested on an NVIDIA 4090D GPU, HybridVLA-dif (7B) and HybridVLA (2.7B) achieve satisfactory control frequencies comparable to CogACT (7B) and $\pi_{0}$ (2.6B), thanks to the reduced DDIM denoising steps and the use of KV cache in HybridVLA. Note that all models are run with bfloat16 precision during inference, without employing action chunking.
4.2 Ablation Study
We conduct ablation experiments on 10 RLBench tasks, using the same training and evaluation settings as in the simulation experiments. To evaluate the effectiveness of the Collaborative Training recipe (CTR), we compare Ex1 with Ex2 and Ex3 with Ex4, as shown in Table 4.2. HybridVLA-dif (Ex1) and HybridVLA-ar (Ex3) are both trained under our proposed CTR that integrates diffusion and autoregressive action generation. Since diffusion tokens precede autoregressive tokens, HybridVLA-dif (Ex1) is evaluated solely on diffusion generation, while HybridVLA-ar (Ex3) performs diffusion denoising followed by autoregressive generation, but is tested only on autoregressive actions. Compared to Ex2 and Ex4, which are trained solely on individual generation methods, both HybridVLA-dif (Ex1) and HybridVLA-ar (Ex3) demonstrate improved manipulation performance. These results validate that our proposed CTR not only avoids negative interference between the two generation paradigms, but also effectively captures the continuous action representations from diffusion-based generation and the pretrained reasoning capabilities from autoregressive generation, enabling mutual reinforcement. The various token formulation designs used in our training recipe are explored in Table 3.2 and Section 3.2. For large-scale pretraining (LSP), we compare Ex5 with Ex0. Although Ex5 is initialized with pretrained VLM parameters, it suffers from a significant drop in accuracy, highlighting the essential role of large-scale pretraining on robot datasets in ensuring stable control. For robot state embedding (RSE), by comparing Ex6 with Ex1, we observe that injecting robot state information enhances the model’s temporal consistency during action prediction. Due to space limitations, Appendix  B.2 provides additional ablation studies on: (1) confidence thresholds in the collaborative action ensemble, (2) the influence of the KV cache on inference speed, and (3) the impact of DDIM sampling steps on performance.
Table 3: Impact of each component. AR and Dif represent autoregressive and diffusion-based action generation, respectively. LSP denotes large-scale pretraining on assembled robotic datasets, while RSE refers to the injected robot state embedding. CTR and CAE represent our proposed collaborative training recipe with hybrid objectives and the collaborative action ensemble method.